Larkspur Corp. had $100,000 of 7%, $20 par value preferred stock and 12,000 shares of $25 par value common stock outstanding throughout 2017.

Assuming that total dividends declared in 2017 were $64,000, and that the preferred stock is not cumulative but is fully participating, common stockholders should receive 2017 dividends of what amount?
Common stockholders should receive $

Assuming that total dividends declared in 2017 were $64,000, and that the preferred stock is fully participating and cumulative with preferred dividends in arrears for 2016, preferred stockholders should receive 2017 dividends totaling what amount?
Preferred stockholders should receive $

Assuming that total dividends declared in 2017 were $30,000, that the preferred stock is cumulative, nonparticipating, and was issued on January 1, 2016, and that $5,000 of preferred dividends were declared and paid in 2016, the common stockholders should receive 2017 dividends totaling what amount?
Common stockholders should receive $
Answer #1

a) Dividends for Preferred Stock = $100,000 * 7% = $7,000

Dividends for common shareholders = $64,000 - $7,000

Dividends for common shareholders = $57,000

b) Dividends for Preferred Stock = $100,000 * 7% * 2 years = $14,000

c) Arrear of Previous year dividend (2016) = 100000* 7% - 5000 = 2000

Current year dividend to preferred stock = 100000*7%= 7000

Total dividend to be paid to preferred stock= 2000+7000= 9000

Dividend to be paid to common shareholders= 30000-9000 = 21000